# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ Abstract
We study the basic properties of resolvent equipped with degenerate metric based on restricted maximal monotonicity, and further propose several conditions for the well-definedness and weak convergence of the fixed point iterations. The results help to understand the behaviours of many operator splitting algorithms in the kernel space of degenerate metric.
@@
Resolvent, degenerate metric, restricted maximal monotonicity, proximal point algorithm (PPA), operator splitting algorithms.
@@ 1.1 Motivation
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued operator. Built on the classical proximal point algorithm (PPA) in 1970s [24, 27], many splitting methods for solving (1) have been developed in the literature during the past five decades. Very recently, a degenerate preconditioned version of PPA was studied in [10], which provided a powerful framework for analyzing various existing splitting methods and also for devising new and flexible schemes. The degenerate preconditioned PPA can be described by the following fixed-point iterate
$x^{k+1}=\mathcal{T}x^{k},\quad\textrm{where\ }\mathcal{T}:=(\mathcal{Q}+\mathcal{A})^{-1}\mathcal{Q},$ (2)
where $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is a linear, bounded, self-ajoint and positive semi-definite operator. To ensure the iterate to be well-defined, the operate $\mathcal{T}$ has to be single-valued and of full domain. However, this key condition was only stated as a basic assumption in [10, Definition 2.1] and hence motivates us to find sufficient guarantees for the well-definedness of $\mathcal{T}$.
Among many operator splitting algorithms prevails the degeneracy phenomenon, which makes the behaviours of (2) more subtle. For instance, the celebrated Douglas–Rachford splitting (DRS), for solving $\min_{u\in\mathcal{U}}(u)+g(u)$, reads as
@@
$\mathcal{H}=\mathcal{U}^{3},\ x^{k}=\begin{bmatrix}u^{k}\\ w^{k}\\ z^{k}\end{bmatrix},\ \mathcal{A}=\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix},\ \mathcal{Q}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&I\end{bmatrix}.$ (4)
Here, we adopt a different (but essentially equivalent) scheme (3) from [10, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate metric $\mathcal{Q}$ in (4) is easier to analyze than that in [10, Eq.(1.5)]. Then, one may ask: Can we conclude the well-definedness of the iterations and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ from (4) only (without ad hoc analysis as [30] did)? More broadly, this further motivates us to study a general setting of (2) and find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$, to guarantee the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$.
@@ 1.2 Related works
When $\mathcal{Q}$ is replaced by a general non-linear operator, the operator $\mathcal{T}$ in (2) was coined the term warped resolvent in [13, Definition 1.1] and was ensured to be well-defined by requiring that $\mathsf{ran}\mathcal{Q}\subseteq\mathsf{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ (see [12, Proposition 2.1] for instance). Besides, a group of sufficient conditions for well-definedness of $\mathcal{T}$ are listed in [13, Proposition 3.9] such as uniform/strict/strong monotonicity. However, these conditions are much strong and hard to verify. Since in this study we only focus on the operator $\mathcal{T}$ with $\mathcal{Q}$ being linear, we are able to find more mild conditions on $\mathcal{A}$ and $\mathcal{Q}$ so that the operator $\mathcal{T}$ to be well-defined.
The convergence issue of degenerate version of (2) has been first tackled in [16, 22] for specific algorithms (corresponding to particular $\mathcal{A}$ and $\mathcal{Q}$) in finite dimensional setting. However, the continuity of the proximity operator—as a main assumption proposed therein—is generally not satisfied in weak topology of infinite dimensional space. For infinite dimensional case, [11] showed the weak convergence for a class of preconditioned ADMM (associated with degenerate preconditioner) based on demiclosedness principle (under additional conditions listed in [11, Lemma 3.4]), [10] further performed a systematic study for general $\mathcal{A}$ and $\mathcal{Q}$. We in this paper for the more general scheme (2) provide more new convergence results with weakened conditions than existing works.
@@ 1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate metric resolvent. The paper is organized as follows. Sect. 2 presents the basic properties of $\mathcal{T}$ under the restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, which are used as preliminary results in Sect. 35 for analyzing full domain, single-valuedness and convergence, respectively. Finally, Sect. 6 exemplifies the results with several splitting algorithms. The contributions are in order.
(1) Full domain: We in Theorem 3.1 give a sufficient and necessary condition for the full domain of $\mathcal{T}$ and further in Corollary 3.7 provide a sufficient condition that is easy to verify in practice. These results are obtained by generalized Moreau’s decomposition identity (cf. Lemma A.1) and restricted Minty’s theorem (cf. Theorem 3.4), based on a notion of restricted maximal monotonicity.
(2) Single-valuedness: Sect. 4 presents sufficient (and necessary) conditions for the single-valuedness of $\mathcal{T}$. Our results weaken the conditions of existing works, from injectivity to restricted injectivity.
(3) Convergence: Theorem 5.3 shows the weak convergence in the range space of $\mathcal{Q}$ under restricted maximal monotonicity only. This result extends [22, Theorem 3.4] to infinite dimensional setting with milder condition than [22]. Notably, the operator $\mathcal{A}$ is not assumed to be maximal monotone for this result. We further show in Theorem 5.6 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details.
@@ 1.4 Notations, assumptions and preliminaries
(cf. [33, Assumption 2.1])
(i) [Basic] $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is linear, bounded, self-adjoint and positive semi-definite, and $\mathsf{dom}\mathcal{Q}=\mathcal{H}$;
(ii) [Degeneracy] $\mathcal{Q}$ is degenerate, i.e., $\mathsf{ker}\mathcal{Q}\backslash\{0\}\neq\varnothing$;
(iii) [Closedness] $\mathcal{Q}$ is closed, i.e., $\overline{\mathsf{ran}\mathcal{Q}}=\mathsf{ran}\mathcal{Q}$.
@@
A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathsf{ran}\mathcal{Q}$ [4, Fact 2.26].
@@
(i) [4, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathsf{ran}\mathcal{Q}+\mathsf{ker}\mathcal{Q}$.
(ii) [4, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x$ are the $\mathsf{ran}\mathcal{Q}$- and $\mathsf{ker}\mathcal{Q}$-parts of $x$.
(iii) [9, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
@@ 2.1 Restricted monotonicity
Under the condition of $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u),(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}).$
The restricted monotonicity is equivalent to the $\mathcal{M}$-monotonicity [10, Proposition 2.5] and also closely related to the local monotonicity [4, Definition 25.7], which requires $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times C)$ to be monotone with the set $C$ being open. The essential difference between the restricted and local versions of monotonicity lies in that $\mathsf{int}\ \mathsf{ran}\mathcal{Q}$ could be empty while $\mathsf{int}\ C$ is always nonempty.
It is well-known that $\mathcal{A}$ is monotone if and only if $\mathcal{A}^{-1}$ is monotone [4, Proposition 20.22]. This fact can be extended directly as below and will provide us with a new perspective for studying the well-definedness of $\mathcal{T}$.
@@
$\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if and only if $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
For convenience of later references, we put Definition 2.1 as
@@
(i) [Basic viability] $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$;
(ii) [Restricted monotonicity] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone.

### v4

# On proximal point method with degenerate preconditioner: well-definedness and new convergence analysis

@@ Abstract
We study the basic properties of degenerate preconditioned resolvent based on restricted maximal monotonicity, and extend the non-expansiveness, demiclosedness and Moreau’s decomposition identity to degenerate setting. Several conditions are further proposed for the well-definedness of the degenerate resolvent and weak convergence of its associated fixed point iterations within either range space or whole space. The results help to understand the behaviours of many operator splitting algorithms, especially in the kernel space of degenerate preconditioner.
@@
Degenerate preconditioner, restricted maximal monotonicity, proximal point method, operator splitting algorithms.
@@ 1.1 Motivation
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued maximally monotone operator. The classical proximal point method (PPM), dating back to 1970s [38, 41], becomes a standard solver of (1), and its close connections to the augmented Lagrangian method (ALM), particularly from the dual viewpoint, have been thoroughly studied in [42, 44, 37].
When the PPM is equipped with a positive definite preconditioner $\mathcal{Q}$, it becomes:
$x^{k+1}=\mathcal{T}x^{k},\quad\textrm{where\ }\mathcal{T}:=(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}.$ (2)
Here, $\mathcal{Q}$ is sometimes called metric [50] and correspondingly, $\mathcal{T}$ is termed preconditioned/metric resolvent [50, 25]. The preconditioned PPM (2) on its own is a conceptual framework, which was shown in [31, 30, 29, 15, 51] to encompass a large variety of first-order operator splitting algorithms, e.g., alternating direction methods of multipliers (ADMM), Chambolle–Pock algorithm [20]. Some extensions of PPM, such as asymmetric forward-backward-adjoint splitting [33, 34], preconditioned forward-backward splitting [52] and TriPD [32], could cover Condat–Vũ [23, 49], and other complicated primal-dual algorithms [16, 11, 12, 24, 21]. $\mathcal{T}$ is a special case of warped resolvent or $\mathcal{Q}$-resolvent of $\mathcal{A}$ [18, Definition 1.1], which allows $\mathcal{Q}$ to be non-linear. (2) is also closely related to the Bregman monotone proximal point algorithm [8, 19, 4], which replaces the standard norm by Bregman distance associated with a Legendre function, and $\mathcal{T}$ is thus also called $D$-resolvent of $\mathcal{A}$ in this context.
Very recently, the linear preconditioner $\mathcal{Q}$ has been extended to be positive semi-definite, rather than strictly positive definite, and thus termed degenerate in [15, 13, 51]. Among many operator splitting algorithms prevails the degeneracy phenomenon. For instance, the celebrated Douglas–Rachford splitting (DRS), for solving $\min_{u\in\mathcal{U}}(u)+g(u)$, reads as
@@
$\mathcal{H}=\mathcal{U}^{3},\ x^{k}=\begin{bmatrix}u^{k}\\ w^{k}\\ z^{k}\end{bmatrix},\ \mathcal{A}=\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix},\ \mathcal{Q}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&I\end{bmatrix},$ (4)
where the preconditioner $\mathcal{Q}$ is clearly degenerate. Here, we adopt a different (but essentially equivalent) scheme (3) from [13, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate preconditioner $\mathcal{Q}$ in (4) is easier to analyze than that in [13, Eq.(1.5)].
The degeneracy of $\mathcal{Q}$ maks the behaviours of (2) more subtle. As mentioned in [13, Remark 2.2], the (global) maximal monotonicity of $\mathcal{A}$ is not sufficient to guarantee the well-definedness of (2). Here the well-definedness of (2) has two-fold implications:
  • $\mathcal{T}$ has to be of full domain, such that (2) admits arbitrary input $x^{k}\in\mathcal{H}$;
  • $\mathcal{T}$ has to be single-valued, such that $x^{k+1}$ can be uniquely determined for the next iterate.
This motivates us to find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$ for the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$. We would also hope, regarding the DRS, to conclude the well-definedness of the iterations (3) and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ using (4) and our general results only (without ad hoc analysis of [47]).
@@ 1.2 Related works
For the general warped resolvent, the well-definedness of $\mathcal{T}$ can be ensured by $\mathrm{ran}\mathcal{Q}\subseteq\mathrm{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ (see [18, Proposition 3.8] and [17, Proposition 2.1] for instance). Besides, a number of sufficient conditions for well-definedness of $\mathcal{T}$ are listed in [18, Proposition 3.9] such as uniform/strict/strong monotonicity. Though these conditions also apply for the degenerate case, they are too strong and hard to verify in practice. More importantly, they failed to associate the maximal monotonicity of $\mathcal{A}$ with degeneracy of $\mathcal{Q}$. The degenerate case is not particularly discussed in [17, 18, 8, 19, 4]. This is our focus to find mild conditions for the degenerate case.
The convergence issue of degenerate version of (2) has been first tackled in [23, 33] for specific algorithms (corresponding to particular $\mathcal{A}$ and $\mathcal{Q}$) in finite dimensional setting. However, the continuity of the proximity operator—as a main assumption proposed therein (cf. [23, Theorem 3.3] and [33, Theorem 3.4])—is generally not satisfied in weak topology of infinite dimensional space. For infinite dimensional case, [15] discussed the weak convergence for a class of preconditioned ADMM, by reformulating it into a degenerate preconditioned PPM, and relying on the demiclosedness (i.e., weak-strong closedness) of $\mathcal{I}-\mathcal{T}$ under several additional assumptions (e.g., compactness and weak closedness in [15, Lemma 3.4]). This analysis was based on the specific structure of $\mathcal{A}$ and $\mathcal{Q}$ (corresponding to ADMM), rather than a general treatment of (2). For the general degenerate setting of (2), [13] performed a systematic study under a basic assumption of the so-called ‘admissible preconditioner[13, Definition 2.1], such that $\mathcal{T}$ is well-defined, which was subsequently used in [27] for studying warped Yosida regularization. In a most recent work of [25], the authors also simply assumed $\mathcal{T}$ to be of full domain (cf. [25, Assumption 1]) and single-valued continuous (cf. [25, Theorems 2.3 and 2.6]). This degenerate preconditioned PPM framework of [13] has been extended or specialized to many following-up variants, e.g., proximal point algorithm with adaptive stepsize [36], fast preconditioned ADMM algorithm [46], distributed DRS based on bilevel graphs [14].
@@ 1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate preconditioned resolvent and its associated iterative scheme (2).
We in Sect. 2.1 introduce a notion of restricted monotonicity, which is much milder than the commonly used (global) monotonicity. Restricted monotonicity connects the monotonicity of $\mathcal{A}$ with the degeneracy of $\mathcal{Q}$. This extends the local monotonicity [5, Definition 25.7], by confining the monotonicity within a subspace instead of an open neighbourhood. This is also somewhat equivalent to the so-called M-monotonicity proposed in [13, Proposition 2.5], which relied on the definition of inversion of the singular preconditioner—$\mathcal{Q}^{-1}$. Instead, we use $\mathcal{A}^{-1}$ in our expositions, and avoid to use $\mathcal{Q}^{-1}$. Lemma 2.9 establishes the $\mathcal{Q}$-democlosedness of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$ for the degenerate case. This is essential for the convergence result (cf. Theorem 5.3). The demiclosedness of $\mathcal{I}-\mathcal{T}$ is rather demanding (see many additional conditions in [15, Lemma 3.4]), we instead propose Lemma 2.9, which, though much weaker than that of $\mathcal{I}-\mathcal{T}$, holds for general $\mathcal{A}$ and $\mathcal{Q}$.
Regarding the full domain of $\mathcal{T}$, Theorem 3.4 presents a sufficient and necessary condition based on a generalized Moreau’s decomposition identity (cf. Lemma 2.11). This is a clear description of ‘admissible preconditioner’ proposed in [13] and adopted by other related works [27, 33, 25]. Furthermore, several counterexamples in Sect. 3.3 show that $\mathcal{T}$ may not have full domain, even if $\mathcal{A}$ is globally maximally monotone, since it cannot guarantee the maximal monotonicity of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$. For the single-valuedness of $\mathcal{T}$, Theorem 4.2 presents several equivalent conditions, which weakens the disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ proposed in [6, 18, 17] to that on $\mathrm{ran}\mathcal{Q}$ with respect to (w.r.t.) $\mathrm{ker}\mathcal{Q}$. This improvement relies to a large extent on the fact that $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}$ is already single-valued (cf. Lemma 2.4).
Theorem 5.3 establishes the weak convergence of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{k}\rightharpoonup\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x^{\star}$ for some $x^{\star}\in\mathrm{zer}\mathcal{A}$, relying on the $\mathcal{Q}$-democlosedness of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$. This extends [33, Theorem 3.4] from infinite dimensional setting to general Hilbert spaces. The convergence within the range space is not considered in [13]. More importantly, Theorem 5.3 does not need to assume the single-valuedness (and hence well-definedness) of $\mathcal{T}$ as [33, Theorem 3.4]. Notably, the operator $\mathcal{A}$ is not assumed to be maximally monotone for this result. Many technical issues are discussed in Sect. 5.2. We further show in Theorem 5.6 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details.
@@ 1.4 Notations, assumptions and preliminaries
(cf. [51, Assumption 2.1])
(i) [Basic] $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is linear, bounded, self-adjoint and positive semi-definite, and $\mathrm{dom}\mathcal{Q}=\mathcal{H}$;
(ii) [Degeneracy] $\mathcal{Q}$ is degenerate, i.e., $\mathrm{ker}\mathcal{Q}\backslash\{0\}\neq\varnothing$;
(iii) [Closedness] $\mathcal{Q}$ has a closed range, i.e., $\overline{\mathrm{ran}\mathcal{Q}}=\mathrm{ran}\mathcal{Q}$.
@@
A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathrm{ran}\mathcal{Q}$ [5, Fact 2.26].
@@
(i) [5, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathrm{ran}\mathcal{Q}+\mathrm{ker}\mathcal{Q}$.
(ii) [5, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathrm{ker}\mathcal{Q}}x$ are the $\mathrm{ran}\mathcal{Q}$- and $\mathrm{ker}\mathcal{Q}$-parts of $x$.
(iii) [10, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
@@ 2.1 Restricted monotonicity
Under the condition of $\mathrm{ran}\mathcal{A}\cap\mathrm{ran}\mathcal{Q}\neq\varnothing$, $\mathcal{A}$ is monotone in ×$\mathcal{H}\times\mathrm{ran}\mathcal{Q}$, or equivalently saying, $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ is monotone [^1] Here, we define a set (rather than an operator) to be monotone. This non-standard but brief terminology is for convenience of our further discussion. See [13, Proposition 2.5] for similar treatment., if
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u),(y,v)\in\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q}).$
The restricted monotonicity is equivalent to the $\mathcal{M}$-monotonicity [13, Proposition 2.5] and also closely related to the local monotonicity [5, Definition 25.7], which requires $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times C)$ to be monotone with the set $C$ being open. The essential difference between the restricted and local versions of monotonicity lies in the fact that for the degenerate case, $\mathrm{ran}\mathcal{Q}$ is not open, and thus, $\mathrm{int}\ \mathrm{ran}\mathcal{Q}=\varnothing$.
It is well-known that $\mathcal{A}$ is monotone if and only if $\mathcal{A}^{-1}$ is monotone [5, Proposition 20.22]. This fact can be extended directly as below and will provide us with a new perspective for studying the well-definedness of $\mathcal{T}$.
@@
$\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ is monotone, if and only if $\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is monotone.
For convenience of later references, we put the assumption of $\mathcal{A}$ as
@@
(i) [Existence of solution to (1)] $0\in\mathrm{ran}\mathcal{A}$, i.e., $\mathrm{zer}\mathcal{A}=\mathcal{A}^{-1}(0)\neq\varnothing$;
(ii) [Restricted monotonicity] $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ is monotone.
Remark 2.3.
For linear operator $\mathcal{Q}$, Assumption 2-(i) also ensures the basic viability of $\mathrm{ran}\mathcal{A}\cap\mathrm{ran}\mathcal{Q}\supseteq\{0\}\neq\varnothing$, which has been mentioned in Definition 2.1, and is a basic viability of $\mathcal{Q}^{-1}\mathcal{A}$ defined in [13, Definition 2.4].